OpenStudy (anonymous):

For breakfast, the cafeteria offers scrambled eggs with one slice of toast or an egg sandwich made with two slices of toast. In fifteen minutes, they sold 16 breakfasts and used 27 slices of bread. Use a model to determine how many egg sandwiches they sold.

OpenStudy (anonymous):

let e be how many scrambled eggs they sold let s be how many sandwiches they sold sounds like e+s = total breakfasts sold while 1*e + 2*s = total slices of bread they needed

OpenStudy (anonymous):

e+s = 16 e+2s = 27 s= 11, e= 5